OpenStudy (anonymous):

how do you change 5% into a fraction lowest terms?

jimthompson5910 (jim_thompson5910):

x% = x/100 In your case, x = 5 So 5% = 5/100 = 1/20

OpenStudy (anonymous):

thanks